Find the ratio of 9m to 27cm.

Grade:12th pass

3 Answers

Pawan Prajapati
2 years ago
In this question, we need to simplify the given ratio in meters and centimeters. For this, we will first make the units of both numbers equal which will be centimeters. Then we will write the ratio in division form and then simplify the division. The obtained division form can be written in ratio form which will be the required ratio. Complete step by step answer: Here, we are given numbers as 9m and 27cm. For writing any two numbers we must have the same units so let us first make their units the same. For this, let us change units of 9 from meters to centimeters. We know that 1 meter is equal to 100 centimeters. Therefore, 9 meter will be equal to 9×100=900 centimeter. Hence, we need to find the ratio of 900cm to 27cm or we can say 900 to 27. Since number before to is written first and number after to is written next. So our ratio looks like this: 900:27 But let us simplify it. For this, let us write it in division form we get: 90027. As we can see, 3 divides both 900 and 27 so dividing the numerator and denominator by 3 we get: 900÷327÷3=3009. Again, 3 divides both 300 and 9, so dividing the numerator and denominator by 3, we get: 300÷39÷3=1003. Our simplified division is 1003. It can be written in ratio as 100:3. Hence 100:3 is our required ratio. Note: Students should always make the units of given numbers the same to calculate ratio. We can convert centimeters to meters also. But it will include decimal which will make it difficult to solve. To check divisibility by 3, we can use the divisibility rule of 3 according to which if the sum of the digits of a number is divisible by 3, then the number itself is divisible by 3.
Aditya Pandey
2 years ago
9m:27cm
1m = 100cm , so 9m = 900 cm
so we get 900:27 = 100:3
so we get that 9m:27cm = 100cm : 3cm
hope it would be helpful to you
Ayush 8A 7239
2 years ago
We first need to convert 9m into cm
8m = 9 × 100 cm = 900cm
ratio of 9m to 27 cm = 900/27  = 300/9 = 100/3 = 100 : 3